# Active particle motion in Poiseuille flow through rectangular channels

### v1

@@ V.3 Motion near equilibrium states
The eigenvectors corresponding to eigenvalues $\lambda_{1,2}=\pm i\sqrt{\bar{U}}$ for the upstream-oriented particle equilibrium are
$\mathbf{v}_{1}=a_{1}\begin{bmatrix}1\\ 0\\ 0\\ -i\sqrt{\bar{U}}\end{bmatrix}+b_{1}\begin{bmatrix}0\\ 1\\ i\sqrt{\bar{U}}\\ 0\end{bmatrix}$
@@
$\mathbf{v}_{2}=a_{2}\begin{bmatrix}1\\ 0\\ 0\\ i\sqrt{\bar{U}}\end{bmatrix}+b_{2}\begin{bmatrix}0\\ 1\\ -i\sqrt{\bar{U}}\\ 0\end{bmatrix}$
respectively. Here, $a_{1},b_{1},a_{2}$ and $b_{2}$ are complex constants with $\bar{a}_{1}=a_{2}$ and $\bar{b}_{1}=b_{2}$, where the overbar denotes a complex conjugate. For small perturbations around this equilibrium point, we numerically observe periodic and quasiperiodic active particle motion confined near the center of the channel (e.g. see Fig. 2(a) and supplemental video S1). We further find that the motion decouples in the $(x,\phi)$ and $(y,\theta)$ variables near this equilibrium point and we obtain the following system for the linearized equation of motion near this equilibrium:
$\ddot{\theta}_{1}+\bar{U}\theta_{1}=0,$
@@
Thus, the evolution of active particle orientations $\theta$ and $\phi$ follow simple harmonic motion with oscillating frequency $\sqrt{\bar{U}}$ and these oscillating orientations drive the translational motion of the active particle near this equilibrium point. Hence, the response to general small perturbations around this equilibrium point is a superposition of the above two decoupled oscillatory motions.
The eigenvectors with eigenvalues $\lambda_{3,4}=\pm\sqrt{\bar{U}}$ for the downstream-oriented equilibrium point are
$\mathbf{v}_{3}=a_{3}\begin{bmatrix}1\\ 0\\ 0\\ \sqrt{\bar{U}}\end{bmatrix}+b_{3}\begin{bmatrix}0\\ 1\\ \sqrt{\bar{U}}\\ 0\end{bmatrix}$
@@
$\mathbf{v}_{4}=a_{4}\begin{bmatrix}1\\ 0\\ 0\\ -\sqrt{\bar{U}}\end{bmatrix}+b_{4}\begin{bmatrix}0\\ 1\\ -\sqrt{\bar{U}}\\ 0\end{bmatrix}$
respectively, with real constants $a_{3},b_{3},a_{4}$ and $b_{4}$. This equilibrium is a saddle point with its unstable manifold tangent to the hyperplane spanned by the two vectors of $\mathbf{v}_{3}$ while its stable manifold is tangent to the hyperplane spanned by the two vectors of $\mathbf{v}_{4}$. To understand the nature of trajectories with small perturbations around this equilibrium point, we simulate motion with different initial perturbations. Some typical trajectories starting near this unstable equilibrium are shown in Fig. 3 with perturbations in the direction of stable/unstable manifolds as well as general perturbation. We see that for a perturbation in the direction of the unstable manifold of the saddle, Fig. 3(a) and supplemental video S8, we obtain a cross-shaped trajectory in the channel cross-section that switches aperiodically between the four diagonals. The chaos here appears to be low-dimensional since the trajectory executes well-defined path on the cross-shape with unpredictability only in the selection of the branches it traverses. For a perturbation in the direction of the stable manifold, Fig. 3(b) and supplemental video S9, we obtain a fan-shaped apparently quasiperiodic trajectory, while for a general perturbation as shown in Fig. 3(c) and supplemental video S10, we obtain a chaotic trajectory with no clear structure. In all cases, we find that small perturbation from this unstable saddle equilibrium leads to wandering-type trajectories that go away from this equilibrium state and explore both the inner and outer regions of the square cross-section.
V.4 Parameter space exploration
Figure 4: (a) Classification of active particle trajectories and (b) largest Lyapunov exponent (LLE) in the initial position $(x(0),y(0))$ plane for the types of trajectories shown in Fig. 2 with fixed $\bar{U}=10$, $\theta(0)=0$ and $\phi(0)=0$. Green ( ) is central swinging motion, cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, yellow ( ) is off-centered trapping motion, purple ( ) is tumbling motion and red ( ) is wandering motion.
We have observed that both the initial conditions and the system parameters can greatly influence the type of active particle trajectory that is realized based on the classification of trajectories described in Fig. 2. In this section, we explore the effects of the initial position and orientation, as well as the parameter $\bar{U}$ on the active particle motion.
@@ V.4.1 Effect of initial position
The active particle dynamics are described by the $4$D nonlinear dynamical system in Eq. (9) which require four initial conditions: two position co-ordinates $x(0)$ and $y(0)$, and two orientation angles $\theta(0)$ and $\phi(0)$. Since there are four initial conditions, we fixed two of them and explore the types of active particle trajectories realized in the initial-condition plane formed by the remaining two initial conditions. We fix the orientation of the active particle to the upstream equilibrium state i.e. $\theta(0)=0$, $\phi(0)=0$, and explore the active particle trajectories in the $(x(0),y(0))$ plane. We consider initial positions restricted to the domain $x,y\in[-0.8,0.8]$. This is done to exclude trajectories that get too close to the wall where interactions of the active particle with the wall may become important. A typical plot at $\bar{U}=10$ depicting the classification of trajectories realized for different initial positions in the cross-section is shown in Fig. 4(a). We find that initial positions near the center of the channel keeps the motion confined near the center of the channel as shown by the green region for central swinging motion (see Fig. 2(a)). An active particle starting out further away from the center of the channel in the vertical and horizontal directions, remains confined in these directions and we observe cyan and blue regions corresponding to vertical swinging and horizontal swinging motion, respectively (see Fig. 2(b)-(c)). Along the diagonals and/or beyond the central region, we have a red region corresponding to wandering trajectories (see Fig. 2(f)). Near the edges of the square cross-section, we have a purple region corresponding to tumbling trajectories (see Fig. 2(e)). Lastly, near the corners, we see islands of yellow regions in the sea of purple corresponding to off-centered confined motion (see Fig. 2(d)).
@@
Figure 5: Trajectory classification (top panels) and LLE (bottom panels) in the initial position $(x(0),y(0))$ plane for fixed $\bar{U}=10$ and varying initial orientations $\theta(0)$ and $\phi(0)$. For fixed $\theta(0)=0$ and (a) $\phi(0)=\pi/24$, (b) $\phi(0)=\pi/12$, (c) $\phi(0)=\pi/6$ and (d) $\phi(0)=\pi/3$. For fixed $\phi(0)=0$ and (e) $\theta(0)=\pi/24$, (f) $\theta(0)=\pi/12$, (g) $\theta(0)=\pi/6$ and (h) $\theta(0)=\pi/3$. Green ( ) is central swinging motion, cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, purple ( ) is tumbling motion and red ( ) is wandering motion. Black regions in panels (d) and (h) correspond to trajectories that came too close to the channel walls i.e. trajectories that went outside a square of $[-0.95,0.95]\times[-0.95,0.95]$.
@@
The above trajectory classification in the initial-condition space is based on the region occupied by the trajectory in the cross-section and does not necessarily capture information about the regular (periodic/quasiperiodic) or chaotic nature of trajectories. However, we typically find that trajectories which are confined near the center of the channel exhibit regular motion whereas active particles that travel near the edges of the cross-section show aperiodic motion and hints of chaos. Furthermore, the trajectories that travel near the edges can flip its sense of rotation in the cross-section between clockwise and counterclockwise (e.g. see tumbling trajectory in supplemental video S5 and S7). The presence of chaos in these trajectories can be quantified by calculating the largest Lyapunov exponent (LLE) of the underlying nonlinear dynamical system [25]. If the LLE is zero, then the active particle motion is either periodic or quasiperiodic, whereas a positive LLE indicates chaos (with the degree of sensitivity to initial conditions given by the magnitude of LLE). We plot the LLE in the plane of initial conditions $(x(0),y(0))$ in Fig 4(b) for the same parameter values as Fig 4(a). We typically find that particles starting in the central region of the square cross-section have regular motion, whereas particles starting near the edges (outer red region and purple region) of the channel are chaotic. However, we also find anomalous periodic regions near corners (within the chaotic sea) where a particle starting in these small regions near the corners of the channel can show regular dynamics. These anomalous periodic regions correspond to the yellow region in Fig. 4(a) indicating off-centered trapping motion.
@@ V.4.2 Effect of initial orientation
We now explore how the classification regions occupied by different types of trajectories in the $(x(0),y(0))$ initial position plane vary with changes in the initial orientation angles $\theta(0)$ and $\phi(0)$ of the active particle around the upstream-oriented equilibrium.
@@
First, we explore the effects of variations in $\phi(0)$ for fixed $\theta(0)=0$. For a small positive value of $\phi(0)=\pi/24$, we find that the blue and cyan regions shrink marginally (see top panel of Fig. 5(a)) compared to the $\phi(0)=0$ case (see Fig. 4(a)). Further increasing to $\phi(0)=\pi/12$, we find that cyan region corresponding to vertically swinging motion vanishes as shown in Fig. 5(b). This is because the value of $\phi(0)$ is no longer close to its upstream equilibrium and hence the larger initial value of $\phi(0)$ are unable to maintain vertical swinging motion of the active particle. Further increasing $\phi(0)$ leads to the shrinkage and disappearance of the green region for central swinging motion too (see Fig. 5(c)-(d)). Again the large value of $\phi(0)$ is unable to maintain a confined trajectory near the channel center. Further, we see that the purple region, corresponding to trajectories confined near the edges, starts mixing with the red region and leads to shrinking of the red region. In terms of the chaotic nature of the trajectories, we see from the bottom panels of Fig. 5(a)-(d), the regular region progressively increases with increasing $\phi(0)$ as a vertical band around the central region of the channel. However, structure indicating chaotic motion also persists within this vertical band of regular motion. The black region near the corners for Fig. 5(d) corresponds to trajectories that were too close to the channel walls i.e. trajectories that went outside the square domain $(x,y)\in[-0.95,0.95]\times[-0.95,0.95]$.
@@
We see a similar trend for variations in $\theta(0)$ for fixed $\phi(0)=0$ with the blue regions replaced with cyan regions. The blue and green regions disappear for large $\theta(0)$ since larger values of $\theta(0)$ cannot give rise to a horizontally swinging motion and trajectories confined to the central region, respectively.
We note that the yellow periodic islands present near the corner of the square cross-section in Fig. 4(a) disappear for variations with initial orientation angle when the other initial orientation angle is kept fixed to zero as shown in Fig. 5. However, similar to cyan/blue regions in Fig. 5(a-d)/(e-h), there does exist a non-zero volume in the initial-condition space where yellow regions would also persist for small variations in $\phi(0)$ and $\theta(0)$ from the upstream equilibrium state.
V.4.3 Effect of $\bar{U}$
@@
Figure 6: Trajectory classification (left panels) and LLE (right panels) in the initial position $(x(0),y(0))$ plane for the types of trajectories shown in Fig. 2 and different values of $\bar{U}$: (a) $\bar{U}=0.5$, (b) $\bar{U}=2.5$, (c) $\bar{U}=5$, (d) $\bar{U}=7.5$ and (e) $\bar{U}=15$. Initial orientation angles are fixed to $\theta(0)=0$ and $\phi(0)=0$. Green ( ) is central swinging motion, cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, yellow ( ) is off-centered trapping motion, purple ( ) is tumbling motion and red ( ) is wandering motion.
@@
Figure 7: Poincaré maps in the $(x_{n},\phi_{n})$ plane for $\theta(t_{n})=0\>\>(\dot{\theta}(t_{n})>0)$ and (a) $H_{s}=1$, (b) $H_{s}=1.5$, (c) $H_{s}=2$, (d) $H_{s}=2.05$, (e) $H_{s}=2.5$ and (f) $H_{s}=3$. The parameter $\bar{U}=10$ was fixed and the initial conditions were chosen based on Eq. (15) to keep $H_{s}$ fixed. Cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, purple ( ) is tumbling motion and red ( ) is wandering motion.
We now look at variations in active particle dynamics with respect to the dimensionless parameter $\bar{U}$ ratio of the maximum flow speed to the intrinsic active particle speed. For $\bar{U}\ll 1$, the active particle intrinsic velocity dominates the flow field and hence, the particle will readily encounter the walls. Moreover, the active particle’s self-generated flow fields would need to be captured to understand its dynamics. Since the simple model used in this work does not capture these effects, we do not explore this regime. This regime of active particle motion in a quiescent fluid within a rectangular duct has been recently explored in detail numerically by Radhakrishnan et al. [26]. Figure 6 shows the different types of trajectories (left panels) and LLE (right panels) in the initial position space for various values of $\bar{U}$ and fixed initial upstream orientation $\theta(0)=0$ and $\phi(0)=0$. For $\bar{U}\lesssim 1$, as shown in Fig. 6(a), we observe regular swinging motion of the active particle about the channel center with the net migration of the particle upstream along the channel. For $\bar{U}\gtrsim 1$, as shown in Fig. 6(b), we again observe regular swinging motion but now the net migration of the particle is downstream along the channel. Increasing to $\bar{U}=5$, the types of trajectories do not change qualitatively compared to Fig. 6(a)-(b), but we start seeing the emergence of chaotic motion for initial conditions near the corners of the channel cross-section as shown in Fig. 6(c). Further increase in $\bar{U}$ to $7.5$ we see the emergence of purple regions near the edges of the cross-section and we also see an increase in the extent of chaotic regions towards the center of the channel (see Fig. 6(d)). For a large value of $\bar{U}=15$, we see the appearance of regular trajectories in the yellow regions near the corners of the channel cross-section. Hence, for the parameter values and the range of $\bar{U}$ values shown in Fig. 6, we find that as $\bar{U}$ increases, the green region corresponding to central swinging motion does not change significantly while the cyan and blue region for vertical and horizontal swinging shrink progressively. Further, with increasing $\bar{U}$, the red region of wandering motion also shrinks with the appearance of tumbling and off-centered trapping motion near the edges and corners of the cross-section. In terms of the chaotic nature, we see a progressive increase in chaotic trajectories up to $\bar{U}=7.5$, but any further increase in $\bar{U}$ leads to the appearance of regions of regular motion near the corners.
@@
Figure 8: Sticky trajectory for active particle in a square cross-section. Particle trajectory in the (a) $(z,y)$ plane and (b) $(x,y)$ plane, as well as (c) Poincaré map in the $(x_{n},\phi_{n})$ plane for $\theta=0\,(\dot{\theta}>0)$. Purple shows the entire trajectory of the particle while light green shows a typical “sticky” region for the trajectory. The parameter $\bar{U}=10$ and initial conditions were chosen such that $H_{s}=3.5$ with $x(0)\approx 0.6861$, $y(0)\approx 0.6582$, $z(0)=0$, $\theta(0)=0$ and $\phi(0)=0$. See also supplemental video S11 for a video of the trajectory.

### v2

@@ V.3 Motion near equilibrium states
The eigenvectors corresponding to eigenvalues $\lambda_{1,2}=\pm i\sqrt{U}$ for the upstream-oriented particle equilibrium are
$\mathbf{v}_{1}=a_{1}\begin{bmatrix}1\\ 0\\ 0\\ -i\sqrt{{U}}\end{bmatrix}+b_{1}\begin{bmatrix}0\\ 1\\ i\sqrt{{U}}\\ 0\end{bmatrix}$
@@
$\mathbf{v}_{2}=a_{2}\begin{bmatrix}1\\ 0\\ 0\\ i\sqrt{{U}}\end{bmatrix}+b_{2}\begin{bmatrix}0\\ 1\\ -i\sqrt{{U}}\\ 0\end{bmatrix},$
respectively. Here, $a_{1},b_{1},a_{2}$ and $b_{2}$ are complex constants with $\bar{a}_{1}=a_{2}$ and $\bar{b}_{1}=b_{2}$, where the overbar denotes the complex conjugate. For small perturbations around this equilibrium point, we numerically observe periodic and quasiperiodic active particle motion confined near the center of the channel (e.g. see Fig. 2(a) and supplemental video S1). We further find that the motion decouples in the $(x,\phi)$ and $(y,\theta)$ variables near this equilibrium point and we obtain the following system for the linearized equations of motion:
$\ddot{\theta}_{1}+{U}\theta_{1}=0,$
@@
Thus, the evolution of active particle orientations $\theta$ and $\phi$ follow simple harmonic motion with oscillating frequency $\sqrt{{U}}$ and these oscillating orientations drive the translational motion of the active particle near this equilibrium point. Hence, the response to general small perturbations around this equilibrium point is a superposition of the above two decoupled oscillatory motions.
The eigenvectors corresponding to eigenvalues $\lambda_{3,4}=\pm\sqrt{{U}}$ for the downstream-oriented equilibrium point are
$\mathbf{v}_{3}=a_{3}\begin{bmatrix}1\\ 0\\ 0\\ \sqrt{{U}}\end{bmatrix}+b_{3}\begin{bmatrix}0\\ 1\\ \sqrt{{U}}\\ 0\end{bmatrix}$
@@
$\mathbf{v}_{4}=a_{4}\begin{bmatrix}1\\ 0\\ 0\\ -\sqrt{{U}}\end{bmatrix}+b_{4}\begin{bmatrix}0\\ 1\\ -\sqrt{{U}}\\ 0\end{bmatrix},$
respectively, with real constants $a_{3},b_{3},a_{4}$ and $b_{4}$. This equilibrium is a saddle point with its unstable manifold tangent to the hyperplane spanned by the two basis vectors defining $\mathbf{v}_{3}$ while its stable manifold is tangent to the hyperplane spanned by the two basis vectors defining $\mathbf{v}_{4}$. To understand the nature of trajectories with small perturbations around this equilibrium point, we simulated motion with different initial perturbations. Some typical trajectories starting near this unstable equilibrium are shown in Fig. 3 for perturbations in the directions of the stable/unstable manifolds as well as a general perturbation. We see that for a perturbation in the direction of the unstable manifold of the saddle, Fig. 3(a) and supplemental video S8, we obtain a cross-shaped trajectory in the channel cross-section that switches aperiodically between the four diagonals. The chaos here appears to be low-dimensional since the trajectory traces a well-defined path on the cross-shape with unpredictability only in the selection of the branches it traverses. For a perturbation in the direction of the stable manifold, Fig. 3(b) and supplemental video S9, we obtain a fan-shaped apparently quasiperiodic trajectory, while for a general perturbation as shown in Fig. 3(c) and supplemental video S10, we obtain a chaotic trajectory with no clear structure. In all cases, we find that a small perturbation from this unstable saddle equilibrium leads to wandering-type trajectories that move away from the equilibrium and explore both the inner and outer regions of the square cross-section.
V.4 Exploration of the parameter space
Figure 4: (a) Classification of active particle trajectories and (b) largest Lyapunov exponent (LLE) in the initial position $(x(0),y(0))$ plane for the types of trajectories shown in Fig. 2 with fixed ${U}=10$, $\theta(0)=0$ and $\phi(0)=0$. Green ( ) is central swinging motion, cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, yellow ( ) is off-centered trapping motion, purple ( ) is tumbling motion and red ( ) is wandering motion.
We have observed that both the initial conditions and the system parameters can greatly influence the type of active particle trajectory that is realized, and have classified these in Fig. 2. In this section, we explore the effects of the initial position and orientation, as well as the parameter ${U}$ on the active particle motion.
The active particle dynamics are described by the $4$D nonlinear dynamical system in Eq. (8) which requires four initial conditions: two position co-ordinates $x(0)$ and $y(0)$, and two orientation angles $\theta(0)$ and $\phi(0)$. To explore the solution space across the four initial conditions, we fix two of them and examine the types of active particle trajectories realized in the initial-condition plane formed by the remaining two.
@@ V.4.1 Effect of initial position
We first fix the initial orientation of the active particle to point upstream, i.e. $\theta(0)=\phi(0)=0$, and explore the variation in active particle trajectories across different initial positions $(x(0),y(0))$. We restrict the domain of initial positions to $x,y\in[-0.8,0.8]$. This is done to exclude trajectories that get too close to the wall where interactions of the active particle with the wall may become important. A plot depicting the classification of trajectories realized for different initial positions in the cross-section when ${U}=10$ is shown in Fig. 4(a). We find that for an initial position near the center of the channel the motion remains confined near the center of the channel as indicated by the green region of Fig. 4(a), and that central swinging motion occurs (see Fig. 2(a)). An active particle starting out further away from the center of the channel but near to an axis remains confined near the same axis as indicated by the cyan and blue regions of of Fig. 4(a), which corresponds to vertical and horizontal swinging motions as in Fig. 2(b,c). Along the diagonals and/or beyond the central region, Fig. 4(a) shows a red region corresponding to wandering trajectories (see Fig. 2(f)). Near the edges of the square cross-section of Fig. 4(a), we have a purple region corresponding to tumbling trajectories (as in Fig. 2(e)). Lastly, near the corners, Fig. 4(a) shows islands of yellow in the sea of purple corresponding to off-centered confined motion (as in Fig. 2(d)).
@@
Figure 5: Trajectory classification (top panels) and LLE (bottom panels) in the initial position $(x(0),y(0))$ plane for fixed ${U}=10$ and varying initial orientations $\theta(0)$ and $\phi(0)$. For fixed $\theta(0)=0$ and (a) $\phi(0)=\pi/24$, (b) $\phi(0)=\pi/12$, (c) $\phi(0)=\pi/6$ and (d) $\phi(0)=\pi/3$. For fixed $\phi(0)=0$ and (e) $\theta(0)=\pi/24$, (f) $\theta(0)=\pi/12$, (g) $\theta(0)=\pi/6$ and (h) $\theta(0)=\pi/3$. Green ( ) is central swinging motion, cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, purple ( ) is tumbling motion and red ( ) is wandering motion. Black regions in panels (d) and (h) correspond to trajectories that came too close to the channel walls i.e. trajectories that went outside a square of $[-0.95,0.95]\times[-0.95,0.95]$.
@@
The above trajectory classification in the initial-condition space is based on the region occupied by the trajectory in the cross-section and does not necessarily capture information about the regular (periodic/quasiperiodic) or chaotic nature of trajectories. However, we typically find that trajectories which are confined near the center of the channel exhibit regular motion whereas active particles that travel near the edges of the cross-section show aperiodic motion and hints of chaos. Furthermore, the trajectories that travel near the edges can flip direction of motion around the origin between clockwise and counterclockwise (e.g. see tumbling trajectory in supplemental videos S5 and S7). The presence of chaos in these trajectories can be quantified by calculating the largest Lyapunov exponent (LLE) of the underlying nonlinear dynamical system [25]. If the LLE is zero, then the active particle motion is either periodic or quasiperiodic, whereas a positive LLE indicates chaos (with the degree of sensitivity to initial conditions given by the magnitude of the LLE). The magnitude of the LLE in the plane of initial conditions $(x(0),y(0))$ is shown in Fig 4(b) for the same domain of parameter values as Fig 4(a). We typically find that particles starting in the central region of the square cross-section have regular motion, whereas particles starting near the walls of the channel (outer red and purple regions of Fig. 4(a)) are chaotic. However, we also find anomalous periodic regions near corners (within the chaotic sea); a particle starting in these small regions shows regular dynamics. These anomalous periodic regions correspond to the yellow region in Fig. 4(a) of off-centered trapping motion.
@@ V.4.2 Effect of initial orientation
We now explore how the classification of trajectories in the $(x(0),y(0))$ initial-position plane varies with small changes in the (fixed) initial orientation angles $\theta(0)$ and $\phi(0)$ from the upstream-orientation.
@@
First, we explore the effect of variations in $\phi(0)$ for fixed $\theta(0)=0$. For a small positive value of $\phi(0)=\pi/24$, we find that the blue and cyan regions shrink marginally from those shown in Fig. 4(a) (where $\phi(0)=0$) to those of the top panel of Fig. 5(a). Further increasing to $\phi(0)=\pi/12$, we find that the cyan region corresponding to vertically swinging motion vanishes as shown in Fig. 5(b). This is because the value of $\phi(0)$ is large enough that, regardless of initial position, the active particle cannot remain confined to the vertical strip which classifies vertical swinging motion. Further increasing $\phi(0)$ to $\pi/6$ and $\pi/3$ leads to the shrinkage and disappearance, respectively, of the green region of central swinging motion (see Fig. 5(c,d)). Again, with a large value of $\phi(0)$ the active particle is unable to remain confined near the channel center. Further, we see that parts of the red region of wandering motion are increasingly replaced by purple tumbling motion with trajectories largely confined near the channel walls. In terms of the chaotic nature of the trajectories, the bottom panels of Fig. 5(a)–(d) show a progressive vertical stretching of the regular region as $\phi(0)$ increases. However, structure indicating chaotic motion also persists within this vertical band of regular motion. The black regions near the corners of Fig. 5(d) correspond to trajectories that came too close to the channel walls, i.e. trajectories that went outside the square domain $(x,y)\in[-0.95,0.95]\times[-0.95,0.95]$. We note that although initial non-zero values of $\phi$ breaks symmetry in $x$ direction, we do largely see a persistent left-right symmetry in Fig. 5. This is probably due to fact that only long-time behavior is captured in the classification of trajectories.
@@
We see a similar trend as $\theta(0)$ is increased for fixed $\phi(0)=0$, but with the blue regions vanishing instead of the cyan regions, see Fig. 5(e–h). The green region also disappears for large $\theta(0)$. Even reasonably small but non-zero values of $\theta(0)$ cannot give rise to a horizontally swinging motion and large $\theta(0)$ also precludes trajectories confined to the central region. Again the presence of up-down symmetry for initial non-zero values of $\theta$ is probably due to fact that only long-time behavior is captured in the classification of trajectories
We note that the yellow periodic islands present near the corner of the square cross-section in Fig. 4(a) are not seen in Fig. 5 when one of the orientation angles is non-zero. Nevertheless, such islands do persist for sufficiently small non-zero values of both orientation angle.
V.4.3 Effect of ${U}$
@@
Figure 6: Trajectory classification (left panels) and LLE (right panels) in the initial position $(x(0),y(0))$ plane for the types of trajectories shown in Fig. 2 and different values of ${U}$: (a) ${U}=0.5$, (b) ${U}=2.5$, (c) ${U}=5$, (d) ${U}=7.5$ and (e) ${U}=15$. Initial orientation angles are fixed to $\theta(0)=0$ and $\phi(0)=0$. Green ( ) is central swinging motion, cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, yellow ( ) is off-centered trapping motion, purple ( ) is tumbling motion and red ( ) is wandering motion.
@@
Figure 7: Poincaré maps in the $(x_{n},\phi_{n})$ plane for $\theta(t_{n})=0\>\>(\dot{\theta}(t_{n})>0)$ and (a) $H_{s}=1$, (b) $H_{s}=1.5$, (c) $H_{s}=2$, (d) $H_{s}=2.05$, (e) $H_{s}=2.5$ and (f) $H_{s}=3$. The parameter ${U}=10$ was fixed and many different initial conditions were chosen based on Eq. (13) while keeping $H_{s}$ fixed. Cyan ( ) is vertical swinging motion, blue ( ) is horizontal swinging motion, purple ( ) is tumbling motion and red ( ) is wandering motion.
We now examine the variations in active particle dynamics with respect to the dimensionless parameter ${U}$, the ratio of the maximum flow speed to the intrinsic active particle speed. For ${U}\ll 1$, the active particle intrinsic velocity dominates the flow field and hence, the particle will readily encounter the walls. Moreover, the active particle’s self-generated flow fields would need to be captured to understand its dynamics. Since the simple model used in this work does not capture these two effects, we do not explore this regime of active particle motion in a quiescent fluid within a rectangular duct. For a recent detailed numerical exploration of this regime see Radhakrishnan et al. [26].
Figure 6 shows the different types of trajectories (left panels) and LLE (right panels) in the initial position space for various values of ${U}$ and fixed initial upstream orientation $\theta(0)=0$ and $\phi(0)=0$. For ${U}\lesssim 1$, and as shown in Fig. 6(a) for $U=0.5$, we observe regular swinging motion of the active particle about the channel center. Moreover, at these low values of $U$ there is a net upstream migration of the particle, i.e. against the flow. For ${U}\gtrsim 1$, and as shown in Fig. 6(b) for $U=2.5$, we again observe regular swinging motion but now $U$ is sufficiently large that the net migration of the particle is downstream in the direction of the flow. Increasing to ${U}=5$, Fig. 6(c), the classification of trajectories does not change qualitatively from that seen at lower values of $U$, but the plot of the LLE shows the emergence of chaotic motion for initial particle positions near the corners of the channel cross-section. Further increasing ${U}$ to $7.5$, Fig. 6(d), we see the emergence of purple regions of tumbling-motion trajectories near the walls of the channel where we also see an increase in the extent of chaotic regions. For a large value of ${U}=15$, Fig. 6(e), we see the appearance of regular trajectories near the corners of the channel cross-section (yellow regions of the trajectory plot).
To summarize, for the parameter values and the range of ${U}$ values shown in Fig. 6, we find that as ${U}$ increases, the green region corresponding to central swinging motion does not change significantly while the cyan and blue regions of vertical and horizontal swinging shrink progressively. Further, with increasing ${U}$, the red region of wandering motion also shrinks with the appearance of tumbling and off-centered trapping motion near the edges and corners of the cross-section. In terms of the chaotic nature, we see a progressive increase in chaotic trajectories up to ${U}=7.5$, but further increase in ${U}$ leads to the appearance of regions of regular motion near the corners.
@@
Figure 8: Sticky trajectory for an active particle in a square cross-section. Particle trajectory in the (a) $(z,y)$ plane and (b) $(x,y)$ plane, as well as (c) Poincaré map in the $(x_{n},\phi_{n})$ plane for $\theta_{n}=0\,(\dot{\theta}_{n}>0)$. Purple shows the entire trajectory of the particle while light green shows a typical “sticky” region for the trajectory. The parameter ${U}=10$ and initial conditions were chosen such that $H_{s}=3.5$ with $x(0)\approx 0.6861$, $y(0)\approx 0.6582$, $z(0)=0$, $\theta(0)=0$ and $\phi(0)=0$. See also supplemental video S11 for a video of the trajectory.